The following are the sample data sets from your class project about waiting line theories:

Arrival: 1, 3, 3, 3, 4

Service time: 33, 33, 52, 69, 58

The arrival data were number of cars arrived per every 5-minute interval, and the service data were the number of seconds spent on taking the corresponding order. What is the expected total number of customers (Cs) in the store for a given 5-minute interval? Give your answer to 2 significant places.
Answer #1

Lambda = average number of arrivals in 5 minutes = 2.8

average service time =49 sec = 49/60 =0.8166 minutes

Mu =average number of persons served in 5 minutes = 5/0.8166 = 6.123

average number of customers in the system = lambda / mu-lambda

= 2.8 / (6.123-2.8) = 0.8426